# On the gravitational stratification of multi-fluid-multi-species plasma

### v1

@@ Abstract
Context. The solar atmosphere is gravitationally stratified and consists of several layers at temperatures different by orders of magnitude. Consequently, the solar atmospheric plasma changes from weakly ionized in the photosphere, partially ionized in the chromosphere, to eventually fully ionized in the corona. However, it is still not trivial to integrate ionization and recombination processes into multi-fluid solar plasma models with gravitational stratification.
@@ 1 Introduction
The solar and stellar atmospheres are highly complex and dynamic. To numerically model and analyze the atmospheres, however, we typically need to make various assumptions and then simplifications. Among them, an important and frequently used one is assuming a 1D ”background” or ”initial” atmosphere in a static gravitational stratification, following
@@
where total pressure includes thermal pressure of gas plasma and magnetic pressure, i.e., $P=p^{\text{g(as)}}+p^{\text{m(agnetic)}}$, $\varrho$ is plasma density, $g$ is gravitational acceleration, and $z$ is height. This balance between the total pressure gradient and gravity is assumed to be valid in the atmospheric model in the $z$-direction. Note that, in the solar atmosphere the mean free path of particles is much smaller than the characteristic length scales of interest. Therefore, the plasma can be assumed to be close to Maxwellian equilibrium, and a fluid description, e.g. magnetohydrodynamics (MHD), is appropriate. Therefore, the stationary momentum equation reduces to the hydrostatic balance as written in Eq. (1). Models with this assumption have been extensively used to investigate, for example, MHD waves in the solar atmosphere, being able to provide important insights (Khomenko & Collados, 2015; Van Doorsselaere et al., 2020).
However, this common practice encounters challenges when considering multi-fluid (two or more fluids) MHD modeling of, e.g., the low solar atmosphere, where the thermodynamic conditions do not permit a single-fluid description, and instead require a set of coupled governing equations for the ionized and neutral components, commonly referred to as multi-fluid MHD equations (Khomenko, 2016; Ballester et al., 2018). While the gravitational acceleration is the same for all fluids, the balance between pressure gradient and density is typically assumed to be independent for each fluid. More specifically, when constructing an initial static atmospheric model, it is frequently assumed that each of the $n$ fluids is in an independent hydrostatic equilibrium
@@
where $N$ and $T$ denote the number densities and temperatures, with subscripts i, n, and e indicating ionized hydrogen, neutral hydrogen, and free electrons, respectively; $k_{\text{B}}$ is the Boltzmann constant. In these models, the description of free electrons is simplified, ignoring the thermal force, which is feasible for modeling the low solar atmosphere which can be assumed to be nearly isothermal (Hansteen et al., 1997). In addition, Eq. (4) assumes that the ion temperature and electron temperature are the same. This assumption simplifies the numerical model but is not a necessary constraint for the solar atmosphere.
@@
Directly using Eq. (2) to calculate a multi-fluid gravitational stratification causes practical issues. Essentially, what Eq. (1) or (2) describes is a boundary value problem that depends on the values of $\varrho_{i}(z_{0})$, in which $z_{0}$ is the reference height, and temperature only plays a role through the equations of state. Therefore, the resulting abundances at certain heights may differ by orders of magnitude from atmospheric models that holistically combine the contributions of necessary physical mechanisms (Vernazza et al., 1981; Fontenla et al., 1993; Avrett & Loeser, 2008). Such differences then change plasma properties and thus the results from numerical models (Zhang et al., 2021). Moreover, as the resulting stratified plasma is largely not in ionization equilibrium, which directly relates to the local temperature, static atmospheric models given by Eq. (2) cannot be used to start simulations with ionization and recombination, which will drastically change the energy balance. Using Eq. (2) also implies the assumption that collisional interactions between different fluids do not affect the static equilibrium. Intuitively, however, one may still expect collisional interactions between fluids, although such interactions may not always be sufficient to fully couple all fluids (Hansteen et al., 1997). Therefore, it is necessary to revise the assumption of independent hydrostatic equilibria for different fluids.
@@
Obtaining an initial atmospheric model that is not only in hydrostatic equilibrium but also in ionization equilibrium, thus far, is not trivial, and only a few two-fluid low solar atmospheric models have considered both equilibria (Maneva et al., 2017; Zhang et al., 2021; Niedziela et al., 2024; Kraśkiewicz & Murawski, 2025). Indeed, it is possible to run a long(er) physical-time simulation to reach a steady state (Brchnelova et al., 2023), in which the initial imbalances disappear. However, the numerical solution depends on the chosen initial conditions, and it is difficult to constrain the solution to obtain a specific atmospheric composition. It can also be computationally intensive. Therefore, to obtain an initial static stratification for a multi-fluid-multi-species (MFMS) MHD model that may include various heavier elements (Wargnier et al., 2023; Martínez-Sykora et al., 2023) in addition to hydrogen, an alternative solution is needed.
@@
To construct an MFMS hydrostatic equilibrium atmospheric model, we assume a static configuration in the absence of external disturbances or force imbalances, such that the characteristic timescale of the equilibrium state is infinitely larger than the collisional timescales. Therefore, we may assume that all the $n$ fluids are coupled by collisional interactions or other mixing processes [^1] This assumption is, however, not valid higher up in the solar atmosphere and the solar wind, where the plasma gradually becomes collisionless and different fluids/species become decoupled (Hansteen et al., 1994; Hansteen & Leer, 1995). The present MFMS model is expected to be used for the low solar atmosphere., approximately acting as a single fluid (Khomenko & Collados, 2012; Martínez-Sykora et al., 2012) in the static state. Then, the gravitational stratification can be calculated straightforwardly. Zhang et al. (2021) used this idea to provide initial equilibria for pure-hydrogen two-fluid plasmas under an isothermal condition. The idea can be extended to MFMS modeling by using a simple numerical integration routine, which is the focus of this work. Below, we start to explain this simple idea with an isothermal example similar to those of Zhang et al. (2021), and then more realistic temperature stratification and ionization are taken into account. We also address the essential physical and numerical consequences of using our model. Additional details are introduced in the Appendix.
@@ 2 An isothermal example
To explain the idea of this work and its necessity, we extend the solution given by Zhang et al. (2021). We further include neutral and ionized helium (He and He+), in addition to the neutral and ionized hydrogen (H and H+), and thus the model has two species and four fluids (not counting free electrons). Extending the solution to more species and fluids is straightforward and will be discussed later.
We assume an upper chromospheric condition adopted from a 2.5D radiative MHD (rMHD) model (Martínez-Sykora et al., 2020a) using the Bifrost code (Gudiksen et al., 2011). The condition was found in regions where the ponderomotive force may produce chemical fractionation, associated with spicules or low-lying loops (Martínez-Sykora et al., 2023). Specifically, the temperature considered is $1.6\times 10^{4}$ K, and the number densities of H, H+, He, and He+ were found to be respectively 109.8, 1010.7, 109.7, and 108.3 per cm3. These number densities are a result of solving for nonequilibrium ionization (NEQ), i.e. solving the full time dependent rate equations for hydrogen and helium. Note that when instead the rMHD model assumes statistical equilibrium (SE) ionization, i.e. solving the rate equations without any $\partial N_{i}/\partial t$ term, where $N_{i}$ is the number density of the $i$-th fluid, the composition includes significantly more neutral fluid. However, in the present example, changing ionization fractions does not affect the conclusions drawn in this paper.
In the original MFMS modeling, gravitational stratification was not considered (Martínez-Sykora et al., 2023). Here, as a first step to improve the MFMS model, we calculate and discuss two different types of stratification:
@@
    Pure hydrostatic equilibrium (pHE): Each fluid is treated independently and satisfies its own hydrostatic equilibrium condition, as given by Eq. 2.
@@
    Coupled hydrostatic equilibrium (cHE): The hydrostatic equilibrium is established through the coupling between different fluids, such that their equilibrium states are mutually dependent.
@@
Figure 1: Density stratifications of hydrogen-helium isothermal mixtures. The solid lines are the cHE stratification and the dashed lines are the pHE stratification, both having the same ionization fractions at the reference height ($z=0$ Mm).
As we assume that in the cHE stratification all fluids are coupled by collisional interactions, at each height all fluids can be described by their bulk motion. Therefore, we define an equivalent atomic mass for considering the bulk motion, which reads
@@
where $m_{i}$ is the atomic mass of the $i$-th fluid. Importantly, free electrons should be considered (thus the nn+1-th fluid) when calculating the equivalent atomic mass, while the electron mass is assumed to be zero. As has been mentioned, in this example we have four ”non-zero-density” fluids and thus $n=4$. Note that we do include separate equations to describe the motion of electrons, as further explained in Section 4. While assuming the ideal gas law, we have the equation of state written as
@@
where $p_{\text{T}}^{\text{g}}$ is the plasma total thermal pressure, $N_{\text{T}}=\sum_{i=1}^{n+1}N_{i}$. Finally, Eq. (1) can be rewritten for the MFMS model, giving us
@@
For the non-magnetic and isothermal case we discuss in this section, an analytical solution
@@
is easily obtained.
In Figure 1, we compare the pHE and cHE results given within a 1 Mm interval [^2] This interval is not suggested to be a realistic layer of the solar atmosphere., using the aforementioned NEQ number densities to provide the plasma quantities at $z=0$, with a constant temperature $1.6\times 10^{4}$ K. The pHE solution results in exponentially changing ionization fractions, especially for neutral helium, whose atomic mass is the largest and requires a larger pressure gradient to balance. In contrast, the cHE solution maintains close to constant ionization fractions for all fluids, which should be the case when the temperature is constant and plasma dynamics are not yet involved. In the cHE solution, each fluid is not in pure hydrostatic balance. Instead, gravity is balanced by the combination of the pressure gradient and the collisional forces, assuming that collisions are sufficiently strong. Therefore, in a cHE stratification the velocity of each fluid is not zero, but their total velocity ($w_{\text{T}}$) along the $z$ direction
@@
is approximately zero. A similar formula is introduced by Khomenko et al. (2014) for the velocity of the center of mass in the one-fluid model, which also explains the static assumption we use before any dynamic driving force is imposed. However, we do not include electrons when defining the total velocity, as the electron velocity is defined as a function of velocities of ions and the total current, i.e. Eq. (13).
Given that ionization fractions affect wave propagation, it is incorrect to use pHE to compute gravitational stratification, except in regions where collisional interactions and other mixing processes are negligible (Hansteen et al., 1997), resulting in decoupled fluids. Moreover, using pHE leads to a higher total density of all fluids, changing wave speeds, etc., thus affecting wave propagation and damping (Zhang et al., 2021). More specifically, the logarithmic slope is different for each of the fluids in pHE vs. cHE. We should also mention that in this simple example the difference in the density distributions of the same fluid in pHE and cHE can already be orders of magnitude. This kind of difference will become more significant when heavier species that have different pressure scale heights are included, and thus it is effectively inviable to use the pHE stratification in MFMS models for evaluating, e.g., the so-called FIP (first ionization potential) effect. Overall, this example shows that it is crucial to construct cHE conditions with physical mixing processes, including, but not limited to, collisional interactions (Hansteen et al., 1994, 1997), particularly for MFMS modeling.
@@ 3 General scenarios
The previous section shows that MFMS plasma can be in a static equilibrium resulting from the balance between gravity, pressure gradients, and collisional forces between fluids. In an isothermal environment, Eq. (7) is easily solved using an analytical solution. In more general scenarios, in which temperature is not constant, ionization fractions should vary accordingly. Therefore, a numerical integration routine is needed.
Below, we assume that the ionization fractions of all species are already given over the entire 1D domain and are not constant. The ionization fractions can be given by assuming the SE condition under realistic solar atmospheric temperature distributions, or provided by self-consistent rMHD simulations. We numerically discretize Eq. (7) to
$\frac{\Delta P_{j}}{\Delta z_{j}}=\frac{\Delta p_{\text{T},j}^{\text{g}}+\Delta p_{j}^{\text{m}}}{\Delta z_{j}}=-\frac{p_{\text{T},j}^{\text{g}}\overline{m}_{\text{T},j}g_{j}}{k_{\text{B}}T_{j}},$ (10)
where the magnetic pressure is given independently. Then we may rewrite Eq. (10) into an integration formula
@@
where $M$ is the number of intervals used to calculate the stratification, and $\overline{m}_{\text{T},j}$ and $T_{j}$ are respectively calculated using Eqs. (5) and (6). $M$ is preferably larger than the number of grid points used in numerical simulations, to reduce the error in the obtained stratification. This is beneficial as the current integration routine is effectively a first-order left Riemann sum, and it can be applied to either a constant or varying $\Delta z$. The present solution is intuitively straightforward, as each integration step is an isothermal scenario introduced in the last section. However, the assumptions we use also imply that neither the hydrostatic equilibrium nor the ionization equilibrium is strictly maintained between any pair of neighboring numerical grid points, thus meaning that a cHE stratification is only in equilibria approximately. Moreover, while collisional interactions are needed to balance gravity and pressure gradient, there must be drift velocities between fluids. Only the total velocity of all fluids is expected to be approximately zero, as has been mentioned above. More details on the remnant disturbances due to the numerical error and the physical decoupling (drift velocities) between fluids are explained in Section 5.

### v2

@@ Abstract
Context. The solar atmosphere is gravitationally stratified and consists of several layers at temperatures different by orders of magnitude. Consequently, the solar atmospheric plasma changes from weakly ionized in the photosphere, partially ionized in the chromosphere, to eventually fully ionized in the corona. However, integrating ionization and recombination processes into multi-fluid solar plasma models with gravitational stratification remains nontrivial.
@@ 1 Introduction
The solar and stellar atmospheres are highly complex and dynamic. To numerically model and analyze the atmospheres, however, we typically need to make various assumptions and simplifications. Among them, an important and frequently used one is assuming a 1D ”background” atmosphere in a static gravitational stratification, following
@@
where total pressure includes thermal pressure of gas plasma and magnetic pressure, i.e., $P=p^{\text{g(as)}}+p^{\text{m(agnetic)}}$, $\varrho$ is plasma density, $g$ is gravitational acceleration, and $z$ is height. This balance between the total pressure gradient and gravity is assumed to hold in the atmospheric model along the $z$-direction. Note that, in the solar atmosphere, the mean free path of particles is much smaller than the characteristic length scales of interest. Therefore, the plasma can be assumed to be in Maxwellian equilibrium, and a fluid description, e.g., magnetohydrodynamics (MHD), is appropriate. Consequently, the stationary momentum equation reduces to the hydrostatic balance as written in Eq. (1). Models with this assumption have been extensively used to investigate, for example, MHD waves in the solar atmosphere, being able to provide important insights (Khomenko & Collados, 2015; Van Doorsselaere et al., 2020).
However, this common practice encounters challenges when considering multi-fluid (two or more fluids) MHD modeling of, e.g., the low solar atmosphere, where the thermodynamic conditions do not permit a single-fluid description, and instead require a set of coupled governing equations for the ionized and neutral components, commonly referred to as multi-fluid MHD equations (Khomenko, 2016; Ballester et al., 2018). In multi-fluid models, the balance between pressure gradient and gravity is typically assumed to be independent for each fluid. More specifically, when constructing an initially static atmospheric model, it is frequently assumed that each of the $n$ fluids is in an independent hydrostatic equilibrium
@@
where $N$ and $T$ denote the number densities and temperatures, with subscripts i, n, and e indicating ionized hydrogen, neutral hydrogen, and free electrons, respectively; $k_{\text{B}}$ is the Boltzmann constant. These two independent equations of state lead to different scale heights of neutral and ionized hydrogen. Moreover, in these models, the description of free electrons is simplified, ignoring the thermal force, which is feasible for modeling the low solar atmosphere which can be assumed to be nearly isothermal (Hansteen et al., 1997). In addition, Eq. (4) assumes that the ion temperature and electron temperature are the same. This assumption simplifies the numerical model but is not a necessary constraint for the solar atmosphere.
@@
Directly using Eq. (2) to calculate a multi-fluid gravitational stratification causes practical issues. As the scale heights are different for different fluids, the resulting abundances at certain heights may differ by orders of magnitude from atmospheric models that holistically combine the contributions of necessary physical mechanisms (Vernazza et al., 1981; Fontenla et al., 1993; Avrett & Loeser, 2008). Such differences then change plasma properties and thus the results from numerical models (Zhang et al., 2021). Moreover, what Eq. (1) or (2) describes is a boundary value problem that depends on the values of $\varrho_{i}(z_{0})$, in which $z_{0}$ is the reference height, and temperature only plays a role through the equations of state. Thus the resulting stratified plasma is largely not in ionization equilibrium, which directly relates to the local temperature, and static atmospheric models given by Eq. (2) cannot be used to start simulations with ionization and recombination, which will drastically change the energy balance. Using Eq. (2) also implies the assumption that collisional interactions between different fluids do not affect the static equilibrium. Intuitively, however, one may still expect collisional interactions between fluids, although such interactions may not always be sufficient to fully couple all fluids (Hansteen et al., 1997). Therefore, it is necessary to revise the assumption of independent hydrostatic equilibria for different fluids.
@@
Obtaining an initial atmospheric model that is not only in hydrostatic equilibrium but also in ionization equilibrium, thus far, is not trivial, and only a few two-fluid low solar atmospheric models (Maneva et al., 2017; Zhang et al., 2021; Niedziela et al., 2024; Kraśkiewicz & Murawski, 2025) have considered both equilibria. Indeed, it is possible to run a long(er) physical-time simulation to reach a steady state (Brchnelova et al., 2023), in which the initial imbalances eventually disappear. However, the numerical solution depends on the chosen initial conditions, and it is difficult to constrain the solution to obtain a specific atmospheric composition. In addition, this strategy can also be computationally intensive. Therefore, to obtain an initial static stratification for a multi-fluid-multi-species (MFMS) MHD model that may include various heavier elements (Wargnier et al., 2023; Martínez-Sykora et al., 2023) in addition to hydrogen, an alternative solution is needed.
@@
To construct an MFMS hydrostatic equilibrium atmospheric model, we assume a static (or steady) configuration in the absence of external disturbances or force imbalances, such that the characteristic timescale of the equilibrium state is infinitely larger than the collisional timescales. Therefore, we may assume that all the $n$ fluids are coupled by collisional interactions (or other mixing processes) [^1] This assumption is, however, not valid higher up in the solar atmosphere and the solar wind, where the plasma gradually becomes collisionless and different fluids/species become decoupled (Hansteen et al., 1994; Hansteen & Leer, 1995). The present MFMS model is expected to be used for the low solar atmosphere., approximately acting as a single fluid (Khomenko & Collados, 2012; Martínez-Sykora et al., 2012) in the static state. Then, the gravitational stratification can be calculated straightforwardly. Zhang et al. (2021) used this idea to provide initial equilibria for pure-hydrogen two-fluid plasmas under an isothermal condition. The idea can be extended to MFMS modeling by using a simple numerical integration routine, which is the focus of this work. Below, we begin by explaining this simple idea using an isothermal example similar to those of Zhang et al. (2021), and then more realistic temperature stratification and ionization are taken into account. We also address the essential physical and numerical consequences of using our model. Additional details are introduced in the Appendix.
@@ 2 An isothermal example
To explain the idea of this work and its necessity, we extend the solution given by Zhang et al. (2021). We further include neutral and ionized helium (He and He+), in addition to neutral and ionized hydrogen (H and H+), so the model has two species and four fluids (not counting free electrons).
We assume an upper chromospheric condition adopted from a 2.5D radiative MHD (rMHD) model (Martínez-Sykora et al., 2020a) using the Bifrost code (Gudiksen et al., 2011). The condition was found in regions where the ponderomotive force may produce chemical fractionation, associated with spicules or low-lying loops (Martínez-Sykora et al., 2023). Specifically, the temperature considered is $1.6\times 10^{4}$ K, and the number densities of H, H+, He, and He+ were found to be respectively 109.8, 1010.7, 109.7, and 108.3 per cm3. These number densities are a result of solving for nonequilibrium ionization (NEQ), i.e., solving the full time-dependent rate equations for hydrogen and helium. When, instead, the rMHD model assumes statistical equilibrium (SE) ionization, i.e., solving the rate equations without any $\partial N_{i}/\partial t$ term, where $N_{i}$ is the number density of the $i$-th fluid, the composition typically includes significantly more neutral fluid.
Based on the given condition, we calculate and discuss two different types of stratification:
@@
    Pure hydrostatic equilibrium (pHE): Each fluid is treated independently and satisfies its own hydrostatic equilibrium condition, as given by Eq. (2).
@@
    Coupled hydrostatic equilibrium (cHE): The hydrostatic equilibrium is established through coupling between different fluids, such that their equilibrium states are mutually dependent.
@@
Figure 1: Density stratifications of hydrogen-helium isothermal mixtures. The cHE stratification and the pHE stratification have the same ionization fractions at the reference height ($z=0$ Mm).
As we assume that in the cHE stratification all fluids are coupled by collisional interactions, at each height all fluids can be described by their bulk motion. We define a mean atomic mass [^2] Hydrogen molecules are largely dissociated at the given temperature. for considering the bulk motion, which reads
@@
where $m_{i}$ is the atomic mass of the $i$-th fluid. We should particularly emphasize that, free electrons should be considered (thus the nn+1-th fluid) when calculating the mean atomic mass, while the electron mass is assumed to be zero. Note that we do include separate equations to describe the motion of electrons, as further explained in Section 4. While assuming the ideal gas law, we have the equivalent equation of state
@@
where $p_{\text{T}}^{\text{g}}$ is the plasma total thermal pressure, and $N_{\text{T}}=\sum_{i=1}^{n+1}N_{i}$ is the total number density. We may also write partial pressures
$p_{i}^{\text{g}}=N_{i}k_{\text{B}}T,\quad i=1,...,n+1,$ (7)
assuming that the temperature is the same for all fluids. Finally, Eq. (1) can be rewritten for the MFMS model, giving us
@@
For the non-magnetic and isothermal ($T\equiv 1.6\times 10^{4}$ K) case here, an analytical solution
@@
is easily obtained. As a result, all fluids/species share the same pressure scale height.
Note that the mean atomic mass is used for calculating the plasma gas pressure. When calculating the (mass) density of a specific fluid, we simply use its number density to multiply the real atomic mass of the fluid. In the present scenario, the ionization fractions do not change over height in the cHE stratification, and thus the (partial) number densities of all fluids can be easily calculated after the total gas pressure has been integrated.
In Figure 1, we compare the pHE and cHE results given within a 1 Mm interval [^3] This interval is not suggested to be a realistic layer of the solar atmosphere., using the aforementioned NEQ condition to provide the plasma quantities at $z=0$. The pHE solution results in exponentially changing ionization fractions, as each fluid has its own scale height, in contrast to the cHE solution. Moreover, in the cHE solution, each fluid is not in pure hydrostatic balance. Instead, gravity is balanced by the combination of the pressure gradient and the collisional forces, assuming that collisions are sufficiently strong. Therefore, in a cHE stratification, the velocity of each fluid is not zero, but their total velocity ($w_{\text{T}}$) along the $z$ direction
@@
is expected to be approximately zero, within the integration error of Eq. (8). A similar form of Eq. (10) is introduced by Khomenko et al. (2014) as the velocity of the center of mass in the one-fluid model, which also explains the static assumption we use before any dynamic driving force is imposed. However, we do not include electrons when defining the total velocity, as the electron velocity is defined as a function of velocities of ions and the total current, i.e., Eq. (13).
Given that ionization fractions affect plasma properties, it is incorrect to use pHE to compute gravitational stratification, except in regions where collisional interactions and other mixing processes are negligible (Hansteen et al., 1997), resulting in decoupled fluids. Having independent scale heights also makes the pHE stratification impractical in MFMS models for evaluating, e.g., the abundances of heavier elements. Overall, this example shows that it is crucial to construct cHE conditions with physical mixing processes, including, but not limited to, collisional interactions (Hansteen et al., 1994, 1997), particularly for MFMS modeling.
@@ 3 General scenarios
3.1 A simple numerical integration routine
In the previous section, we used a simple example to analytically construct a coupled MFMS stratification that is in balance between gravity, pressure gradients, and collisional forces between multiple fluids/species. In more general scenarios, in which temperature is not constant, ionization fractions should vary even if SE is assumed. Therefore, a numerical routine is needed to integrate Eq. (8).
Below, we assume that the ionization fractions of all species are already calculated over the entire 1D domain based on any given temperature distribution, and are not constant. The ionization fractions can be given by assuming the SE condition under realistic solar atmospheric temperature distributions, or provided by self-consistent rMHD simulations with, for instance, NEQ ionization. We then discretize Eq. (8) and obtain a numerical integration formula
@@ 3.1 A simple numerical integration routine
where $M$ is the number of intervals, with $\Delta z_{j}=z_{j+1}-z_{j}$, $\overline{m}_{\text{T},j}$ is calculated using Eq. (5), and $T_{j}$ is given and not changed by the integration routine. $\overline{m}_{\text{T},j}$ and $T_{j}$ are also considered to be constant in each integration step. More specifically, at discrete point $j$, the local ionization fraction is used to calculate local partial number densities, which are then used in Eq. (5). Except at $j=0$, where the boundary solution is imposed, the local total number density is calculated using (I) the total gas pressure provided by the integration step j1j-1, and (II) the equivalent equation of state, Eq. (6). Note that the contribution of magnetic pressure is discussed in Appendix A.
In practice, $M$ is preferably larger than the number of grid points used in numerical simulations, to reduce the error in the obtained stratification. This is beneficial as the present integration routine is effectively a first-order left Riemann sum. The present solution is intuitively straightforward, as each integration step is an isothermal scenario introduced in the last section. However, the assumptions we use also mean that neither the hydrostatic equilibrium nor the ionization equilibrium is strictly maintained between any pair of neighboring numerical grid points due to numerical error. Moreover, while collisional interactions are needed to balance gravity and pressure gradient, there must be drift velocities between fluids. Only the total velocity of all fluids is expected to be approximately zero, as has been mentioned above.